Thanks for the previous answers. I really need help to this maths; it's that urgent.
Which of the following is not a zero of
P(x) = -5(x - 14)(x + 10)(x - 11)(x - 4)(x + 16)?
Option: 5, 4, 11, -10
Please write the solution procedure for me.

Answer by checkley77(12844)

OPTION 5 IS THE ONLY ONE THAT DOESN'T TURN ONE OF THE (TERMS)=0
PROOF:
(-10+10)=0
(4-4)=0
(11-11)=0